OpenStudy (anonymous):

(2x − 3)^3 Perform operations and simplify

Nnesha (nnesha):

are you working on binomial theorem ?

Nnesha (nnesha):

binomial theorem for (x-y)^3\[\huge\rm (x-y)^3 = x^3-3x^2y+3xy^2-y^3\]

OpenStudy (anonymous):

can you tell me the answer please i've already tried the question and got the worng ansewr @Nnesha

Nnesha (nnesha):

alright then show ur work plz so i can find out the mistakes so we don;t have to restart it again

OpenStudy (anonymous):

i don'thave me work with me...it's an online assignment...

Nnesha (nnesha):

no i mean how did you get the wrong answer ? i just want to know where u did a mistake

OpenStudy (anonymous):

I don't know where i made a mistake...

Nnesha (nnesha):

\[\huge\rm (x-y)^3 = x^3-3x^2y+3xy^2-y^3\]use this equation cube of first term `-` 3 times (square of first term times 2nd term) + 3 times first term times square of 2nd term - cube of 2nd term

Nnesha (nnesha):

ye ik so that's why i'm saying post your work here maybe i will be able to find the mistake :=)

OpenStudy (anonymous):

I'm not goingto type all my work out right now

Nnesha (nnesha):

ok.as you wish .

Nnesha (nnesha):

i asked a question would you like to answer that or no sir ????

OpenStudy (anonymous):

First off...I'm not a sir...Secondly it's not a quiz or anything, i got the answer wrong multiple times and would just like to know the answer...

Nnesha (nnesha):

alright sorry about that mam i can't give you the answer you got the answer wrong it's okay i'm saying it would be great to just look over ur work and find the mistake instead wasting time doing it again.

Nnesha (nnesha):

by just looking at the answer how would you find the mistake ?

Nnesha (nnesha):

anyways i'll do the first term (x-y)^3\[\huge\rm (x-y)^3 =\color{ReD}{ x^3}-\color{blue}{3x^2y}+3xy^2-y^3\] in your question 2x is the first term \[\huge\rm (2x-y)^3= \color{Red}{(2x)^3 - \color{blue}{3(2x)^2(-y)}}\]

OpenStudy (anonymous):

I've done it with multiple problems...if i know the answer i can get the work easier

Nnesha (nnesha):

better use calculator then. sorry